### v2

# FLUXSynID: A Framework for Identity-Controlled Synthetic Face Generation with Document and Live Images

Abstract
Synthetic face datasets are increasingly used to overcome the limitations of real-world biometric data, including privacy concerns, demographic imbalance, and high collection costs. However, many existing methods lack fine-grained control over identity attributes and fail to produce paired, identity-consistent images under structured capture conditions. We introduce FLUXSynID, a framework for generating high-resolution synthetic face datasets with user-defined identity attribute distributions and paired document-style and trusted live capture images. The dataset generated using the FLUXSynID framework shows improved alignment with real-world identity distributions and greater inter-set diversity compared to prior work. The FLUXSynID framework for generating custom datasets, along with a dataset of 14,889 synthetic identities, is publicly released [^1] https://github.com/Raul2718/FLUXSynID to support biometric research, including face recognition and morphing attack detection.
1 Introduction
Large-scale, diverse, and well-annotated face datasets are crucial for training and evaluating biometric systems. In addition to face recognition [40, 44, 49] and emotion analysis [51, 15], recent work on morphing attack detection (MAD) [7, 34, 2] and face demorphing [35, 53] also depends heavily on such datasets. However, collecting real-world facial data presents substantial challenges. Privacy regulations restrict access and usage, while the high financial cost of acquiring and annotating large volumes of data can be prohibitive. Moreover, demographic imbalances and the need for operationally relevant conditions, such as capturing both document-style and trusted live images for the same subject, further complicate data collection efforts.
While generative models such as Variational Autoencoders (VAEs) [30], GANs [20], and Diffusion Models [50] have enabled synthetic face dataset creation [39, 5, 45, 6, 27], most existing methods focus on in-the-wild imagery and lack support for structured capture conditions or fine-grained control over identity attributes (e.g., age, ethnicity). Moreover, generated faces often diverge from the distributional characteristics of real data [25]. The ONOT dataset [14] addresses some of these issues but remains limited in scale, control, and intra-set diversity.
In this work, we introduce FLUXSynID, a framework for generating synthetic face datasets with controllable identity attributes and paired document-style and live capture images. Built on the FLUX.1 [dev] diffusion model [3] and enhanced via LoRA-based fine-tuning [24], FLUXSynID produces high-quality, identity-consistent images guided by natural language prompts. It employs a two-stage pipeline: generating document-style images first, then corresponding live capture variants using three complementary methods. These are based on LivePortrait [21], PuLID [22], and Arc2Face [42], each introducing different levels of variation in pose, expression, lighting, and background.
The contributions of this work are threefold:
  • A method for generating high-quality document-style face images with user-defined identity attribute distributions, allowing direct demographic control.
  • A multi-method approach to generating synthetic trusted live capture images that preserve identity while introducing realistic capture variability.
  • A comparative evaluation of FLUXSynID against comparable real (FRGC [43], FRLL [11]) and synthetic (ONOT [14], StyleGAN2 [27]) face datasets, demonstrating improved alignment in visual appearance and in embedding space distributions produced by face recognition systems.
To support research in morphing attack detection and related biometric fields, we publicly release the FLUXSynID framework along with a dataset of 14,889 synthetic identities.
2 Related Work
Synthetic face datasets are increasingly used in biometric research due to the limitations of real facial data, such as privacy issues, high costs, and demographic imbalances [25, 39, 14, 5]. Early methods such as SynFace [45] and DigiFace-1M [1] showed promise but lacked realism and flexibility. SynFace used a 3D-aware GAN to generate facial images with limited variation, while DigiFace-1M used 3D rendering to create large-scale identities but struggled to match real-world distributions and offered limited control over attributes such as age. Both fell short in visual fidelity and attribute precision, limiting their use in tasks such as morphing attack detection (MAD).
GANDiffFace [39] is a recent approach combining two generative models to achieve explicit control over demographics and identity. It uses StyleGAN3 [26] to generate high-resolution synthetic faces and applies linear Support Vector Machines (SVMs) for targeted manipulation of demographic attributes. To introduce intra-class variation, a diffusion model is fine-tuned using DreamBooth [52] on each generated identity, allowing for controlled sampling. Despite these advancements, GANDiffFace, similar to SFace [5], focuses on unconstrained images and lacks document-style outputs. StyleGAN3, while powerful, struggles with consistent identity preservation across variations and requires a separate SVM for each attribute. Furthermore, the need to fine-tune a diffusion model per identity adds significant computational overhead, limiting the scalability of this approach for biometric tasks.
Recent studies have leveraged diffusion models for identity-consistent face generation, achieving greater visual realism and intra-class variation. DCFace [29] introduced a dual-conditioning approach using both identity and style embeddings to generate varied images of the same person. IDiff-Face [4] extended this by conditioning a latent diffusion model on face embeddings from recognition networks, producing distinct yet consistent identities. Both demonstrate strong recognition performance and show that diffusion models better capture real-world identity spaces. However, they focus on unconstrained synthesis and do not support document-style imagery paired with live captures. Additionally, identity attribute control is limited to what the embeddings encode, lacking mechanisms for explicit demographic customization.
The ONOT dataset [14] is the only prior work focused on document-style face synthesis for biometric use. It employed a fine-tuned Stable Diffusion [50] model with prompt engineering to generate ICAO-compliant [56] identity images, sampling 64 images per identity across 15k identities and retaining only those with at least one compliant image. However, this brute-force filtering resulted in usable samples for just 27% of identities, with further intra- and inter-class filtering reducing the dataset to a maximum of 255 identities. Our analysis shows that even these retained identities were highly similar and poorly aligned with the real identity embedding space. ONOT’s document-style control relies solely on prompts, lacking the fine-tuning necessary for consistent ID-photo outputs. Its demographic attribute control is limited and not easily extensible to new or composite features. While ONOT shares our high-level goals, it falls short in flexibility, scalability, and representational diversity.
3 FLUXSynID: Document Images
Figure 1: Overview of the FLUXSynID pipeline for synthetic document image generation. Given user-defined identity attributes and constraints, descriptive prompts are generated and used to guide FLUX.1 [dev] [3] diffusion model enhanced with LoRA adapter [24] for producing realistic frontal ID-like images.
Figure 1 presents an overview of the framework for the generation of synthetic document images in FLUXSynID dataset. The FLUX.1 [dev] [3] diffusion model is used to generate high-quality, high-resolution images based on user-defined identity attributes.
3.1 Identity Attributes
Depending on the application, existing face image datasets may exhibit imbalances in identity attributes such as gender, age, or ethnicity [25]. To address this, the FLUXSynID framework allows user-defined attribute classes, each with a set of attributes with associated selection probabilities. Additionally, each class can be assigned an inclusion probability during sample generation. As shown in Figure 1, the gender class is always included (100%), with the Female identity attribute selected 49% of the time. Other classes, such as body type, may be included in only a portion of samples (e.g., 50%), enabling fine-grained control over attribute distribution in the synthetic dataset.
However, some attribute combinations can lead to semantic inconsistencies. For example, assigning Bald hairstyle along with a hair color such as Black creates a contradiction. To address this, FLUXSynID allows the definition of attribute clashes. Users can specify primary and secondary attributes, where the primary is retained and conflicting secondaries are discarded. In the above example, if Bald hairstyle is selected, any hair color attribute is excluded to ensure consistency.
Once a set of attributes is selected for a synthetic identity, it is stored for future reference, and a Large Language Model (LLM) then generates a coherent textual prompt for image synthesis. The use of an LLM is motivated by the fact that FLUX.1 [dev] [3] employs a T5 [47] encoder, which excels at capturing nuanced semantic relationships from verbose, detailed prompts. Specifically, the pre-trained Qwen2.5 [57] model is employed to generate natural language descriptions that consistently and semantically integrate the selected attributes.
3.2 Document Image LoRA Adapter
Low-Rank Adaptation (LoRA) [24] enables efficient fine-tuning of large pre-trained models by introducing a small set of trainable parameters. Instead of updating the full model weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the LoRA adapter keeps $W_{0}$ fixed and learns a low-rank update $\Delta W=BA$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$. The adapted weights become $W=W_{0}+\Delta W$. This approach significantly reduces trainable parameters compared to methods such as DreamBooth [52], used in generating the GANDiffFace [39] dataset.
In this work, a LoRA adapter is used to fine-tune the FLUX.1 [dev] [3] diffusion model for stylistic consistency in image generation. The goal is to generate high-resolution, vivid frontal face images with neutral expressions and a consistent white background. The adapter is trained on images from the Chicago Face Database (CFD) [37, 38, 32], which consists of high-quality portraits that match these criteria. Figure 2 shows outputs with and without LoRA, highlighting improved pose, lighting, and expression consistency.
Figure 2: Comparison of synthetic document images generated with and without LoRA [24]. Without LoRA, images exhibit inconsistent head poses, facial expressions, and lighting conditions. LoRA fine-tuning produces more uniform, ID-style images with neutral expressions and consistent photographic conditions.
3.3 Image Generation
Once trained, the LoRA adapter is used to update the weights of the pre-trained and frozen FLUX.1 [dev] [3] model, which conditions image generation based on prompts produced by Qwen2.5 [57]. FLUX.1 [dev] is a rectifying flow transformer [16] latent diffusion [50] model tailored for text-to-image synthesis. As a guidance-distill model, FLUX.1 [dev] incorporates Classifier-Free Guidance (CFG) [23] directly into the transformer through a modulation scale referred to as a guidance scale, eliminating the need for negative prompts and halving inference steps by avoiding separate conditional and unconditional passes. The guidance scale directly controls prompt adherence: lower values result in more diverse outputs with weaker attribute fidelity, while higher values tend to produce oversaturated, less varied images.
As shown in Figure 1, the generated prompt $P$ is encoded using the pre-trained and frozen CLIP-L [46] and T5-XXL [47] models. FLUX.1 [dev] leverages these encoders complementarily: CLIP’s global embedding is combined with the timestep embedding and used to modulate transformer layers, while T5-XXL’s token embeddings are concatenated with latent image tokens to form a unified sequence processed via self-attention. This achieves effects similar to cross-attention, allowing text tokens to guide specific image regions [19].
After encoding the text prompt, $P$, using both encoders ($E_{\text{CLIP-L \& T5-XXL}}$), the resulting textual conditioning, $c_{P}$, is used as conditional input to FLUX.1 [dev] alongside a 16-channel Gaussian noise latent image $z_{t}$ sampled from a random seed. Through latent diffusion, FLUX.1 [dev] iteratively denoises $z_{t}$ over $t$ steps to produce the final latent image. The conditional denoising process is defined as:
$z_{t-1}=\text{DPM++ 2M}\left(z_{t},t,\epsilon_{\theta^{CFD}}\left(z_{t},t,c_{P}\right)\right),$ (1)
where DPM++ 2M [36] with the beta schedule [33] is used as the sampling algorithm in diffusion probabilistic modeling, and $\epsilon_{\theta^{CFD}}$ represents the FLUX.1 [dev] model with the LoRA adapter trained on CFD images (see Section 3.2).
The final denoised latent image, $z_{0}$ is then passed to the pre-trained FLUX.1 [dev] Variational Autoencoder (VAE) decoder network [3], $D_{\text{VAE}}$, which generates the synthetic document image, $I_{D}$, containing the identity attributes described in the textual prompt.
4 FLUXSynID: Trusted Live Capture Images
Figure 3: Overview of the FLUXSynID live capture image generation pipeline. Starting from a synthetic document image, three methods are used to produce identity-consistent live images: LivePortrait [21] enables controlled variations in pose and expression; Arc2Face [42] introduces natural diversity via identity embeddings; and the PuLID-based approach [22, 54] is integrated with FLUX.1 [dev] [3] via identity and edge-based conditioning, guided by modified image prompts.
Figure 3 shows the framework for generating synthetic trusted live capture images for the FLUXSynID dataset. Three methods are used: LivePortrait [21], Arc2Face [42], and PuLID [22], producing live images $I_{LL}$, $I_{LA}$, and $I_{LP}$, respectively. Each method provides unique strengths in identity preservation, expression, lighting, and pose variation, making them well-suited for diverse applications.
4.1 LivePortrait-based Image Generation
LivePortrait [21] is a facial animation method that modifies facial expressions and head orientations from a single input image. Unlike diffusion-based methods, it employs an implicit-keypoint approach that allows for greater efficiency and fine-grained control. Its retargeting mechanism employs compact keypoints to enable subtle adjustments to facial features and pose, with a stitching module that ensures visual coherence by seamlessly blending altered facial regions.
Within FLUXSynID, LivePortrait generates modified versions of document images, introducing controlled variations randomly sampled for each image. This preserves identity while subtly adjusting pose and expression.
4.2 Arc2Face-based Image Generation
Arc2Face [42] is an identity-conditioned image generation framework that synthesizes realistic facial images using identity embeddings extracted via ArcFace [12]. It adapts a pre-trained Stable Diffusion [50] model to generate images directly from identity vectors, ensuring strong identity preservation. Unlike LivePortrait [21], Arc2Face introduces natural variations in pose, expression, and lighting without explicit user control. The authors report superior identity similarity compared to other generative models, underscoring its effectiveness. Notably, Arc2Face requires no textual prompts, relying entirely on image-based identity embeddings.
In FLUXSynID, Arc2Face is used to generate synthetic images directly from ArcFace embeddings of the original synthetic document images. This introduces natural diversity in expressions, poses, and lighting while maintaining identity consistency. The method complements the more controlled LivePortrait approach, enriching the dataset with greater variability.
4.3 PuLID-based Image Generation
PuLID (Pure and Lightning ID Customization) [22] personalizes text-to-image diffusion models using identity-specific embeddings derived from reference images. Unlike traditional fine-tuning methods, PuLID allows for rapid identity injection without the need for retraining each subject. It uses a dual-path training strategy: one path conditioned on text only, and the other on both text and identity embeddings, with contrastive and identity loss to align outputs.
Within FLUXSynID, we employ PuLID-FLUX-v0.9.1 [54], an adaptation of PuLID integrated with the pre-trained and frozen FLUX.1 [dev] diffusion model [3]. As shown in Figure 3, the synthetic document image $I_{D}$ and the LivePortrait-modified image $I_{LL}$ are passed through PuLID’s identity encoder. The resulting embeddings are averaged to form the identity condition $c_{ID}$, capturing both static identity features and dynamic pose or expression cues.
PuLID introduces dedicated cross-attention layers into FLUX.1 [dev]’s transformer blocks to integrate $c_{ID}$. Unlike text embeddings, which are fused via global modulation and token concatenation, identity embeddings interact directly with the FLUX.1 [dev] transformer layers via these specialized attention modules.
Simultaneously, the original identity prompt $P$ is transformed into a contextual prompt $P^{\prime}$, describing the same individual in an airport setting, and encoded via CLIP-L [46] and T5-XXL [47] to produce the textual condition $c_{P^{\prime}}$.
In parallel, edges from $I_{LL}$ are extracted and encoded into a latent representation $z_{t,\text{edges}}$ using the pre-trained FLUX.1 [dev] VAE encoder ($E_{\text{VAE}}$[3]. Simultaneously, a Gaussian noise latent image $z_{t,\text{noise}}$ is generated with the same seed as the document image. Both latent images are concatenated across the channel dimension to form $z_{t}^{*}$:
$z_{t}^{*}=z_{t,\text{edges}}\mathbin{\|}z_{t,\text{noise}}.$ (2)
The concatenated latent image $z_{t}^{*}$ serves as the input for the FLUX.1 [dev] diffusion model equipped with a specialized FLUX.1 [dev] Canny LoRA adapter [3]. This LoRA adapter modifies the input layer of FLUX.1 [dev] and is explicitly trained for edge-based image-to-image conditioning.
Therefore, the final generation is conditioned on three signals: identity embedding ($c_{ID}$), modified text embedding ($c_{P^{\prime}}$), and latent edge representation ($z_{t,\text{edges}}$), enabling realistic, identity-consistent live capture synthesis. The conditional denoising process in PuLID-based live image generation is formulated as:
$z_{t-1}^{*}=\text{DPM++ 2M}\left(z_{t}^{*},t,\epsilon_{\theta^{\text{Canny}}}\left(z_{t}^{*},t,c_{P^{\prime}},c_{\text{ID}}\right)\right),$ (3)
where $\epsilon_{\theta^{Canny}}$ represents FLUX.1 [dev] model combined with the FLUX.1 [dev] Canny LoRA adapter [3].
5 Experiments
5.1 Implementation Details
Prompt Generation.
Identity prompts were generated using the Qwen2.5 [57] model, based on sampled identity attributes. Attribute inclusion probabilities and conflict rules were applied to generate semantically rich prompts optimized for conditioning FLUX.1 [dev] [3].
LoRA Training for Document Image Generation.
A LoRA adapter was trained using FluxGym [9] on 830 images from the Chicago Face Database [37, 38, 32], with rank $r=16$, learning rate 0.0002, and 18 training epochs. Image captions were generated using the InternVL2.5 [8] vision-language model. The LoRA module was activated during inference via a unique token (“s7ll2f4h”), and additional parameters in the CLIP-L [46] encoder were also fine-tuned.
Document Image Generation.
Document images were synthesized through the ComfyUI-based [10] implementation of FLUX.1 [dev] [3], conditioned on Qwen2.5 prompts. Guidance scales were sampled randomly between 1.7 and 2.5 to balance fidelity with identity diversity. Each image was generated using $t=20$ diffusion steps.
Live Image Generation.
LivePortrait [21] and PuLID-FLUX-v0.9.1 [54] were used via modified ComfyUI-based [10] implementations. A fixed guidance scale of 4 and $t=20$ steps were used with PuLID, as recommended by the authors. Arc2Face [42] was run via the official implementation with default hyperparameters.
Dataset Composition.
A total of 15,000 synthetic identities were generated across 14 attribute classes: gender, age, region of origin, body type, eye shape, lip shape, nose shape, face shape, hairstyle, hair color, eyewear, facial hair, skin type, and ICAO-compliant headwear [56]. The gender, age, and region of origin attributes were included in 100% of prompts to guarantee consistent representation. The probabilities for age, region of origin, and body type attributes were uniformly distributed, whereas other attributes were subjectively assigned, reflecting author preference instead of real-world distributions. Researchers are encouraged to define custom distributions for task-specific dataset generation.
Of the 15,000 identities, 111 were excluded due to Arc2Face [42] failing to extract identity embeddings. These identities were removed to ensure each identity includes outputs from all three live image generation methods.
All images were generated at a resolution of $1024\times 1024$, except for Arc2Face-based live images, which were constrained to $512\times 512$ due to implementation limitations.
5.2 Similarity-Based Identity Filtering
Due to prompt similarity and inherent biases in the FLUX.1 [dev] [3] model, some synthetic identities may appear overly similar. Unlike ONOT [14], which enforces strict identity separation by removing all similar identities detected by a face recognition system (FRS), FLUXSynID allows a small degree of similarity, reflecting real-world datasets where a limited False Match Rate (FMR) is acceptable. For instance, FRONTEX guidelines [18] permit an FMR of 0.1%. As dataset size grows, the probability that a randomly selected identity $i$ falsely matches at least one other identity in the dataset is:
$P_{\text{False Match}}(i)=1-(1-\text{FMR})^{N-1},$ (4)
where $N$ represents the total number of identities.
To manage this behavior, FLUXSynID employs iterative similarity-based filtering. Starting from an initial set of document images $D_{o}$, pairs of identities with cosine similarity exceeding a threshold are identified. The filtering process iteratively removes identities involved in the highest number of such matches. This continues until the dataset-wide FMR reaches or falls below a predefined static FMR target:
$\text{FMR}_{D_{f}}\leq\text{FMR}_{\text{target}}.$ (5)
Here, dataset-wide FMR is calculated as the ratio of the number of pairs exceeding the threshold to the total number of possible comparisons. This method ensures the final subset $D_{f}\subseteq D_{o}$ maintains a practical and acceptable FMR.
Experiments in this study leverage two open-source face recognition models: ArcFace [12] and AdaFace [28]. For each model, two decision thresholds are selected, corresponding to FMR targets of 0.1% and 0.01%. These thresholds are derived from 340k impostor verification attempts on identities from the CFD [37, 38, 32] dataset, yielding values of 0.423 and 0.497 for ArcFace, and 0.253 and 0.334 for AdaFace, respectively. Accordingly, the $\text{FMR}_{\text{target}}$ in Equation (5) is set to 0.1% or 0.01%, depending on the chosen threshold.
The number of identities remaining after filtering varies significantly depending on the FRS used and the FMR target. Table 1 summarizes the results across both ArcFace and AdaFace for target FMRs of 0.1% and 0.01%.
FRS FMR Target Remaining Identities (%)
ArcFace 0.01% 9,358 (62.9%)
ArcFace 0.1% 6,641 (44.6%)
AdaFace 0.01% 6,074 (40.8%)
AdaFace 0.1% 2,389 (16.0%)
Table 1: Remaining identities after similarity-based filtering at different FMR targets using ArcFace [12] and AdaFace [28]. Percentages are relative to the initial 14,889 identities.
For comparison, we also applied ONOT’s [14] stricter identity filtering, which removes all identity pairs exceeding the threshold, effectively enforcing a dataset-wide FMR of zero. Using ArcFace with ONOT’s specified thresholds, FLUXSynID retained 51.0% of identities at FMR 0.1% and 87.1% at FMR 0.01%. In contrast, ONOT reports retaining only 3.1% and 6.3%, respectively. Although the comparison is not exact due to ONOT’s use of additional intra-class filtering, these results suggest that FLUXSynID identities exhibit substantially greater feature-level diversity.
5.3 Evaluation
Visual Results.
Figure 4 shows sample identities from FLUXSynID, each consisting of a document image ($I_{D}$) and live capture images generated by LivePortrait ($I_{LL}$[21], PuLID ($I_{LP}$[22], and Arc2Face ($I_{LA}$[42]. Document images are frontal, feature neutral expressions, and have uniform white backgrounds, reflecting stylistic control via CFD-based LoRA fine-tuning.
Each live image generation method introduces different levels of variation. LivePortrait introduces subtle, controlled changes in pose and expression while maintaining consistent lighting and background, although minor blurring is sometimes evident. PuLID leverages text, edge maps, and identity embeddings to alter both background and pose, although it may occasionally cause identity drift. Arc2Face generates the most diverse outputs, introducing substantial changes in setting, lighting, and expression that closely mimic in-the-wild photographs.
Figure 4: Examples of synthetic identities from FLUXSynID, each shown with a document image ($I_{D}$) and corresponding live capture images generated using LivePortrait ($I_{LL}$[21], PuLID ($I_{LP}$[22], and Arc2Face ($I_{LA}$[42].
Prompt-Image Consistency.
Figure 5 shows synthetic document images generated from randomly sampled identity attributes. Generally, the images align well with the specified attributes, although some limitations are evident. Age estimation occasionally deviates, producing individuals who appear significantly younger or older than intended, consistent with findings in [41], which highlight FLUX.1 [dev] [3]’s difficulty in accurately representing age.
Figure 5: Examples of randomly sampled identity attributes and their corresponding synthetic document images ($I_{D}$) generated using FLUXSynID framework.
Subtle facial features, such as nose shape, are also inconsistently realized. This may stem from limitations in automated image captioning, which often omits fine-grained details when annotating datasets for training text-conditioned diffusion models [31, 13, 17].
While increasing the guidance scale improves prompt adherence, excessively high values can introduce artifacts and reduce realism. Empirically, guidance scales between 1.7 and 2.5 yielded the best trade-off between fidelity and prompt consistency.
Impact of Similarity-Based Filtering on Attribute Distribution.
To evaluate the potential bias introduced by similarity-based filtering (Section 5.2), Figure 6 compares the distributions of gender, age, and region of origin attributes before and after filtering.
Female identities were observed to be removed more frequently, reducing their overall representation. This indicates that FLUX.1 [dev] [3] generates less diverse female faces, leading to higher similarity scores and more frequent exclusion. Conversely, male and non-binary identities increased proportionally, indicating greater internal variability.
In the age category, younger individuals were more frequently filtered out, reflecting reduced variation among these groups. Attribute distributions for region of origin remained largely stable, although some shifts were observed. For example, there was a decline in Central American identities and an increase in East African ones, highlighting uneven feature diversity across regions that may stem from biases in the generative model.
Figure 6: Distribution of identity attributes in the dataset before and after filtering based on identity similarity. Each bar plot shows the percentage of images with a given attribute. Blue bars represent the original dataset, and orange bars represent the dataset after filtering.
Distribution of Identity Embeddings.
Figure 7 visualizes identity embeddings extracted by ArcFace [12] and AdaFace [28] across real and synthetic datasets. Synthetic StyleGAN2 [27] images were generated using the pSp [48] encoder, producing random identities in frontal poses. For ONOT [14], ICAO-compliant images from Subset 1 (the largest identity set) were used. FLUXSynID identities were filtered using FMR@0.01% thresholds (Section 5.2) to remove highly similar identities. Only document images were considered to maintain consistency across datasets, all of which contain document-like image types.
The results show that StyleGAN2 and ONOT identities form distinct clusters that do not overlap well with real-world identity distributions from CFD [37, 38, 32] and FRLL [11]. In contrast, FLUXSynID embeddings demonstrate greater overlap with real data, suggesting closer alignment with real-world identity variability. Imperfect overlap may reflect FLUXSynID’s broader demographic range, particularly in age and region of origin, compared to the more limited scope of CFD and FRLL.
Moreover, ONOT identities form tight clusters, indicating low inter-class diversity. In contrast, FLUXSynID embeddings are more evenly distributed, capturing a broader and more realistic spectrum of identity features.
(a) ArcFace [12]
(b) AdaFace [28]
Figure 7: t-SNE [55] visualization of FRS embeddings derived from real and synthetic identity images.
Similarity Score Distributions.
(a) FLUXSynID-LL
(b) FLUXSynID-LA
(c) FLUXSynID-LP
(d) FRLL [11]
(e) FRGC [43]
(f) ONOT [14]
Figure 8: Distribution of cosine similarity scores for mated (blue) and non-mated (orange) identity pairs across real and synthetic datasets, using the AdaFace FRS with an FMR@0.01% threshold.
Dataset Mated Non-Mated
FRLL FRGC FRLL FRGC
ONOT 23.793 6.770 5.202 4.715
FLUXSynID-LL 1.047 15.512 0.248 0.375
FLUXSynID-LP 23.992 1.655 0.332 0.477
FLUXSynID-LA 19.225 0.145 0.038 0.101
Table 2: KL divergence between similarity score distributions of synthetic and real datasets for mated and non-mated identity pairs. Lower values indicate greater distributional similarity.
To further evaluate FLUXSynID, cosine similarity scores for mated pairs (document image and its corresponding live image) and non-mated pairs (document image compared to live images from 100 different identities) are visualized as histograms in Figure 8. Comparisons are made to real-world FRLL [11] and FRGC [43] datasets.
Results indicate that FLUXSynID-LL (LivePortrait) closely matches FRLL, reflecting high identity consistency with minimal appearance variation. FLUXSynID-LP (PuLID) and FLUXSynID-LA (Arc2Face) better align with FRGC, reflecting broader visual diversity. ONOT exhibits overlapping mated and non-mated distributions that are skewed relative to real data, indicating limited identity separation and low inter-class variability.
These observations are quantified in Table 2, which reports KL divergence between synthetic and real datasets. FLUXSynID-LL yields the lowest mated-pair divergence from FRLL, supporting its close visual match and controlled variation. For non-mated pairs, FLUXSynID-LA achieves the lowest divergence, though its mated-pair scores are higher. Overall, FLUXSynID-LL offers the most balanced similarity to FRLL across both pair types.
For FRGC, FLUXSynID-LA records the lowest mated-pair KL divergence, although FLUXSynID-LP (PuLID) visually resembles FRGC images more closely. While its KL values are slightly higher than Arc2Face’s, they still indicate strong distributional similarity.
In contrast, ONOT exhibits high KL divergence across all comparisons, indicating that its similarity score distributions differ substantially from those of real datasets. This observation aligns with Figure 8, where ONOT’s mated and non-mated distributions are visibly shifted compared to real data.
6 Conclusion
We present FLUXSynID, a synthetic face dataset generation framework capable of producing high-resolution, identity-consistent document and live capture images with fine-grained control over identity attributes. Combining prompt-driven diffusion synthesis with LoRA fine-tuning [24] and three complementary live image generation methods, FLUXSynID facilitates both controlled and naturalistic variations across identity instances. Experimental results demonstrate that FLUXSynID aligns more closely with the distributional characteristics of real-world datasets than existing synthetic alternatives, while its embedding space diversity supports realistic biometric evaluation scenarios.
Despite its strengths, FLUXSynID does have some limitations. Attributes such as age and subtle facial features are not always represented reliably, and similarity-based filtering may unintentionally skew demographic distributions. In contrast to ONOT, which enforces strict ICAO compliance [56] at the cost of a 73% reduction in dataset size [14], FLUXSynID does not enforce ICAO constraints. This decision reflects the limitations of current compliance tools, which frequently involve opaque checks and can result in compounding errors. Instead, FLUXSynID prioritizes dataset size and embedding space diversity over strict ICAO adherence, enhancing its flexibility for a broader range of biometric research applications.
7 Acknowledgments
This research was funded by the European Union under the Horizon Europe programme, Grant Agreement No. 101121280 (EINSTEIN).